2012 CERN Beam Test Video

CERN BT Report, Dec. 3rd 2012

Position scan with CREAM-2 HPD boxes in configuration 1 (front beam without lead) was completed.
Position scan with CREAM-2 HPD boxes in configuration 2 was processed until 8 am on Dec 3.
At 8 am on Dec 3 (Monday), beam stopped.
The instrument on the rotation fixture was moved from beam line to testing area.
Targets were unmounted from the pallet and stand-off under the target legs were removed.
TCD was removed. Targets were re-mounted with shims.
SBT and rotation fixture controller were uninstalled.
Client PC and (ethernet and serial) cables were uninstalled.
Packing started.

JiHye worked on removing LPSC's HPDs from ISS-CREAM HPD boxes
in order to disassemble an electronics board for HPD mounting on each HPD box.
Those boards were requested to be packed to UMD by MooHyun.
With the help of David, Mike, HyunGue and Kenny, wires on HPDs were cut with longest length possible, 
HPDs were un-mounted from HPD boxes, and the electronics boards were dis-assembled from the HPD box.
Electronics boards were packed into the crate.
LPSC's HPDs were put into a storage box one by one carefully.
Ludo picked up HPDs and HPD boxes to take LPSC on Dec. 4, Tuesday.
(When JiHye was trying to start working on HPD boxes, Jason pointed out that
she should follow the ESD training contents.
Since there was not ESD work bench and we did not bring ESD suit nor wrist strip,
we have to proceed it carefully. At least the table was grounded on the wall.)

Sasha helped with the KNU shipping arrangement to Korea.

 

-YS

 

TAMU Radiation Test Report, May 28th, 2012

We have completed the testing of 6 chips at around 4:45 AM (CDT, Local time) which were scheduled to be completed by May 31st 8 AM thanks to the newly designed two test boards which have three chips per each. It reduced the beam time by a factor of three. So instead of 32 hours, we could finish in about 10 hours.

Here is a test summary.

1. ADC chip: Start to show SEL (Single Event Latch-up) at around 10.4 MeV cm2/mg with Argon particles of 15 MeV/nucl. This corresponding to about 10-3/cm2/day and with 1000 days of flight, we will see one SEL per chip during the duration. And for the 80 ADC chips in the CAL, it means about one SEL per 12 days. So this chip can be used in the CREAM.

2. DAC chip: No SEL up to 49.3 MeV cm2/mg with Xenon particles. 10-5 SEL for 1000 days per chip.

3. 74HC123 (Monostable vibrator): No SEL up to 49.3 MeV cm2/mg. 10-5 SEL for 1000 days per chip.

4. MUX chip: The same as the above two. Shows current increase at LET of 85.4 MeV cm2/mg with Gold particles but still working.

5. Comparator chip: Single event upset from LET of 22.8 MeV cm2/mg but no SEL.

6. Current monitor chip: No SEL/SEU up to 49.3 MeV cm2/mg.

In conclusion, all 6 chips are tested and found to be OK in 1000 days of flight in the ISS.

 

Best, MooHyun